# Efficient Biological Data Acquisition through Inference Set Design

### v2

@@ Abstract
In drug discovery, highly automated high-throughput laboratories are used to screen a large number of compounds in search of effective drugs. These experiments are expensive, so one might hope to reduce their cost by experimenting on a subset of the compounds, and predicting the outcomes of the remaining experiments. In this work, we model this scenario as a sequential subset selection problem: we aim to select the smallest set of candidates in order to achieve some desired level of accuracy for the system as a whole. Our key observation is that, if there is heterogeneity in the difficulty of the prediction problem across the input space, selectively obtaining the labels for the hardest examples in the acquisition pool will leave only the relatively easy examples to remain in the inference set, leading to better overall system performance. We call this mechanism inference set design, and propose the use of an confidence-based active learning solution to prune out these challenging examples. Our algorithm includes an explicit stopping criterion that stops running the experiments when it is sufficiently confident that the system has reached the target performance. Our empirical studies on image and molecular datasets, as well as a real-world large-scale biological assay, show that active learning for inference set design leads to significant reduction in experimental cost while retaining high system performance.
@@ 1 Introduction
Automated high-throughput screening (HTS) laboratories have enabled scientists to screen large compound libraries to find effective therapeutic compounds and screen whole-genome CRISPR knockouts to understand the effects of genes on cell function (Mayr & Bojanic, 2009; Wildey et al., 2017; Blay et al., 2020; Tom et al., 2024; Fay et al., 2023). However, conducting experiments on every compound or gene in these vast design spaces remains very resource-intensive. With typical screening libraries holding on the order of $10^{5}$ to $10^{6}$ compounds (Hughes et al., 2011) and the number of possible small molecules estimated at $10^{60}$ (Bohacek et al., 1996), the disparity between our screening capabilities and all which we could explore is staggering. Reducing experimental costs without compromising the quality of the generated data would allow us to accelerate biology and pharmaceutical research and expand the set of molecules considered for testing.
@@
To avoid costs scaling with the number of experiments, we can train a model from a subset of the target library that has been tested in the lab, and then predict experimental outcomes for the remainder of the library using the trained model (Naik et al., 2013; Reker & Schneider, 2015; Dara et al., 2022), thereby building a hybrid screen of the library. This approach entails three interrelated questions: (1) which subset of the library should we use to maximise the accuracy of the predictions?, (2) how do we select this subset without access to the experimental outcomes?, and (3) how do we ensure that we acquire a large enough collection to meet a target level of accuracy for the predictions?
@@
This problem is similar to an active learning problem in that we want to select examples that maximize prediction accuracy, but instead of aiming to minimize generalization error, here we focus solely on prediction on a particular, finite, set of experiments from a library. The fact that the library is finite introduces an important difference: the learner can influence the set of examples on which it is evaluated by strategically selecting examples to be acquired. We assume that the difficulty of predicting outcomes varies among examples. Some experiments outcomes are inherently harder to predict either due to their complex properties (Bengio et al., 2009), because of the partial observability of the system (Saleh et al., 2021; Krenn et al., 2020) or due to noise in the labeling function (Frénay & Verleysen, 2013; Lukasik et al., 2020). Such disparity allows the learner to select these harder examples into the training set to avoid having to predict their outcomes. Conversely, if an example can be reliably predicted by the model, we can save experimental costs by not including it in the training set. We call this mechanism through which a learner can influence evaluation performance inference set design.
@@
We propose an active learning-based solution to hybrid screening that uses the model’s confidence to guide the selection of experiments and leverages the mechanism of inference set design to improve the system’s performance on the target set. Our algorithm includes a practical stopping criterion that terminates the acquisition of additional labels once a lower bound on a target accuracy threshold is exceeded, and show that this bound provides probabilistic guarantee on the performance of the algorithm as long as the model is weakly calibrated. To validate our method, we conduct a series of empirical studies on image and molecular datasets, as well as a real-world case study in drug discovery. The results demonstrate that inference set design significantly reduces experimental costs while improving overall system performance. Importantly, this is true even when the generalization benefits of active learning-based acquisition functions are marginal compared to random search. This has important practical implications for active learning: if a problem is a hybrid screen—in the sense that one only needs good performance on a fixed, finite set of experiments—then evaluating generalization error dramatically understates the benefits of active learning. By combining simple active learning acquisition functions with an appropriate stopping criterion, it is possible to make large scale screening far more efficient.
@@ 2 Methods
In this section, we present the problem of efficiently evaluating a large, finite set of experimental designs, such as a screening compound library. We formalize the problem as a subset selection problem, and show that it can be solved using an active learning strategy. Active learning is effective in this setting because it essentially selects its own inference set; we call this strategy, inference set design. We we propose a stopping criterion to monitor the model’s performance and trigger termination of the acquisition process and show that this algorithm allows to reach provably high levels of performance on a target set of samples. A pseudo-code is available in Appendix B.
@@ 2.1 Reducing Experimental Costs via Hybrid Screens and Inference Set Design
We are motivated by the problem of efficient data acquisition of compound libraries for drug discovery applications. Each “experimental readout” (or label) $y$ requires us to run an experiment, which has some associated cost. We assume that we have a predefined target set $\mathcal{X}_{\text{target}}=\{x_{i}\}^{N_{\text{target}}}_{i=1}$ of experimental designs for which we want to acquire a readout, where $N_{\text{target}}$ represents the number of samples we wish to obtain labels or predictions for. For example, in drug discovery, a common set of experimental designs, $\mathcal{X}_{\text{target}}$, would be a library of compounds to be tested on some set of cell types, at different concentrations, etc. For each $x_{i}\in\mathcal{X}_{\text{target}}$, the associated experimental readout, $y_{i}$, would correspond to the drug’s effect on the corresponding cell type. In this section, we assume a classification problem but also explore extensions to continuous outcomes in Appendix E.1. One way to reduce the acquisition costs of screening the entire target set is to train a predictive model $\hat{p}(x)$ on the observation set – the subset of the target set, $\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}_{\text{target}}$, for which we already have observed the labels $y$. We can then use this model’s predictions in place of the real labels on the remaining samples. Because experiments are typically run in multiple rounds, at each round $t$, the target set $\mathcal{X}_{\text{target}}$ can be partitioned into two mutually exclusive subsets:
@@
  • The observation set $\mathcal{X}_{\text{obs}}^{t}$, which contains the union of all experimental designs that have already been tested in the lab and for which the readouts, $\mathcal{Y}_{\text{obs}}^{t}$, have been observed. We denote their pairing the observation dataset $\mathcal{D}_{\text{obs}}^{t}:=(\mathcal{X},\mathcal{Y})_{\text{obs}}^{t}:=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{obs}}}$. At each step $t$, a constant number of $N_{b}$ samples are acquired and added to the observation set, causing this set to grow over time.
@@
  • The inference set $\mathcal{X}_{\text{inf}}^{t}:=\mathcal{X}_{\text{target}}\setminus\mathcal{X}_{\text{obs}}^{t}$, which consists of the remaining experimental designs in the target set that have not yet been tested. This set shrinks over time as samples are selected from it for acquisition and transfered into the observation set.
@@
Once acquisition is stopped, at timestep $t=\tau$, the output of the system is a hybrid dataset, i.e. the combination $(\mathcal{X},\mathcal{Y})_{\text{obs}}^{\tau}\cup(\mathcal{X},\hat{\mathcal{Y}})^{\tau}_{\text{inf}}$ of both the labeled pairs from the observation set and the predicted pairs from the inference set. We call this procedure a hybrid screen, as it provides a hybrid set of labels for the entire target set – some measured, some predicted. The system’s performance at each step $t$ is evaluated based on its accuracy $\mu_{\text{sys}}^{t}$ across the entire target set, which consists of perfectly accurate labels $y$ from the observation set and the inferred labels $\hat{y}$ for the inference set:
@@
where $|\mathcal{X}_{\text{obs}}|$ denotes the number of samples in the observation set. A natural approach is to treat this as an active learning problem. Like standard active learning, we can sequentially select subsets of data from labeling and the system accuracy will benefit from any improvements in generalization that arise from better selection of training examples. However, this setting has an important difference: we only ever evaluate $\hat{p}$ on $\mathcal{X}_{\text{inf}}$, and this set is under the control of the learner because any example that is selected for labeling is removed from the inference set. As a result, by choosing acquisition functions that select harder training examples firsts, we make the test-time task of the learner easier, as the harder examples are effectively pruned out of the inference set. We call this approach inference set design – by selective label acquisition, the agent actively designs the composition of the inference set such that it can excel on it. This mechanism is illustrated in Figure 1.
@@ 2.2 Empirical Stopping Criterion
The key technical challenge of inference set design is deciding when to stop running new experiments. In a production system, the labels for the inference set will remain unavailable, so we need to estimate the system performance from the collected data only. To avoid unnecessary experimentation, inference set design algorithms must employ an efficient stopping criterion. We address this by maintaining a probabilistic lower bound on the system accuracy, and stopping once this bound exceeds the critical threshold, $\gamma$. A simple approach to maintaining such a bound is to leverage the feedback from each round of experimentation.
@@
$x_{\text{selected}}:=\operatorname*{arg\,max}_{x_{i}\in\mathcal{X}_{\text{inf}}\setminus\mathcal{X}_{b}}\Big{(}1-\hat{v}_{i}\Big{)}\quad\text{for}\quad i=1,\dots,N_{b}$ (3)
@@ 3 Related Work
In this work, we apply active learning to the problem of hybrid screens for biological data acquisition. Warmuth et al. (2003) were among the firsts to apply AL to the field of drug discovery, with an acquisition function based on support vector machines for binding affinity predictors. AL has since then been explored as a tool for problems such as virtual screening (Fujiwara et al., 2008), cancer monitoring (Danziger et al., 2009), protein-protein interaction (Mohamed et al., 2010), compound classification (Lang et al., 2016) and chemical space exploration (Smith et al., 2018).
These approaches seek to produce a better model than training on random samples would, and are to be distinguished from bayesian optimization methods (Graff et al., 2021; Gorantla et al., 2024), which also fall under the umbrella of “active learning” but instead use the model to seek specific samples (e.g. with high binding affinity) and often focus on addressing the exploration-exploitation dilemma which has been widely studied by bandit algorithms (Svensson et al., 2022). For hybrid screens, we seek a model that can accurately predict the labels of all samples in the inference set, without preference over particular readouts. Our setting, focusing on a fixed set of data, is sometimes referred to as pool-based active learning (Wu, 2019; Zhan et al., 2021). A key difference is that most active learning work (including pool-based AL) tend to focus evaluations on a held-out test set (Atlas et al., 1989; Cohn et al., 1994; Gal et al., 2017; Margatina et al., 2021; Zhan et al., 2021; Luo et al., 2023; Li et al., 2024) or restrict the algorithm to very limited acquisition budgets (e.g. only $10\%$ of the data pool) (Wu, 2019), which generally does not leave enough margin for mechanisms such as inference set design to dominate.
@@ 4 Experiments
In our experiments, we aim at highlighting the key role that inference set design plays into allowing AL agents to reach high levels of performance on the inference set (Section 4.1), validating the robustness of this mechanism on molecular datasets used for drug discovery (Section 4.2), applying this method to a challenging real-world case of compound library screening on a large scale phenotypic assays (Section 4.3) and empirically validating our theoretical assumption and stopping criterion (Section 4.4). All curves show average metrics across 3 seeds and the shaded area denote standard error around the mean. Vertical lines denote the average stopping time, with min and max intervals (left to right). For all modalities, we use vectorized representations as input and our model $\hat{p}$ is parameterized using a feed-forward deep neural network with residual connections. Additional details regarding data pre-processing, network architectures and hyperparameters are presented in Appendices C and D. The code for all experiments will be made public upon publication of this work. All datasets to reproduce our results are also publicly available, except one proprietary dataset for the results in Figure 8.
Figure 2: Performance of least-confidence (LC) and random acquisition functions on variations of MNIST across 3 seeds. a) Original: MNIST with its naturally ocurring easy and hard examples. b) Partial Observability: MNIST where the bottom two thirds of the images have been cropped out. c) Noisy labeling function: MNIST with shuffled labels for 6, 8, and 9 digits.
Figure 3: Performance of active and random agents on MNIST with increasing number of shuffled labels. As the task becomes harder, the accuracy gap increases, but the stopping time $\tau$ moves back.

### v3

@@ Abstract
In drug discovery, highly automated high-throughput laboratories are used to screen a large number of compounds in search of effective drugs. These experiments are expensive, so one might hope to reduce their cost by only experimenting on a subset of the compounds, and predicting the outcomes of the remaining experiments. In this work, we model this scenario as a sequential subset selection problem: we aim to select the smallest set of candidates in order to achieve some desired level of accuracy for the system as a whole. Our key observation is that, if there is heterogeneity in the difficulty of the prediction problem across the input space, selectively obtaining the labels for the hardest examples in the acquisition pool will leave only the relatively easy examples to remain in the inference set, leading to better overall system performance. We call this mechanism inference set design, and propose the use of a confidence-based active learning solution to prune out these challenging examples. Our algorithm includes an explicit stopping criterion that interrupts the acquisition loop when it is sufficiently confident that the system has reached the target performance. Our empirical studies on image and molecular datasets, as well as a real-world large-scale biological assay, show that active learning for inference set design leads to significant reduction in experimental cost while retaining high system performance.
@@ 1 Introduction
High-throughput screening (HTS) laboratories have enabled scientists to efficiently perform whole-genome CRISPR knockouts and screen large compound libraries to discover effective therapeutics (Mayr & Bojanic, 2009; Wildey et al., 2017; Blay et al., 2020; Tom et al., 2024; Fay et al., 2023). However, conducting such experiments on every compound or gene in these vast design spaces remains resource-intensive. With typical screening libraries holding on the order of $10^{5}$ to $10^{6}$ compounds (Hughes et al., 2011) and the number of possible small molecules estimated at $10^{60}$ (Bohacek et al., 1996), the disparity between our screening capabilities and all which we could explore is staggering. Reducing experimental costs without compromising the quality of the generated data would allow to greatly accelerate biological and pharmaceutical research.
@@
At any stage of a screening procedure, to reduce experimental costs, we can train a machine learning model on the data that has already been collected, and then use that model to predict experimental outcomes for the remainder of the experiments (Naik et al., 2013; Reker & Schneider, 2015; Dara et al., 2022). This procedure generates a hybrid screen of the target library, with some outcomes experimentally observed and others obtained from model predictions. However this approach entails interrelated questions: Which subset of the library should we use to maximize the accuracy of the predictions? How do we select this subset without access to all experimental outcomes? How do we ensure that we acquire a large enough collection to meet our accuracy goal?
@@
This setting is similar to an active learning problem in that we want to select examples that maximize prediction accuracy, but instead of aiming to minimize generalization error on the entire data space, we focus solely on prediction on one particular set of experiments. The fact that this library is finite introduces an important difference: here the learner can influence the set of examples on which it is evaluated by strategically selecting the examples to be acquired. Since it is often the case that some experiments outcomes are inherently harder to predict than others, either due to their complex properties (Bengio et al., 2009), because of the partial observability of the system (Saleh et al., 2021; Krenn et al., 2020) or due to noise in the labeling function (Frénay & Verleysen, 2013; Lukasik et al., 2020), the learner can select these harder examples into the training set early on to avoid having to predict their outcomes at a later stage. Conversely, if a group of examples can be reliably predicted by the model, we can save experimental costs by not including them in the training set. We call this mechanism through which a learner can influence its own evaluation set: inference set design.
@@
In this work, we propose an active learning-based solution to hybrid screening. Our approach uses the model’s confidence to guide the selection of experiments and leverages the mechanism of inference set design to improve the system’s performance on the target set. Our algorithm includes a practical stopping criterion that terminates the acquisition of additional labels once a lower bound on a target accuracy is exceeded, and we show that this bound provides probabilistic guarantee on the performance of the algorithm as long as the model is weakly calibrated. To validate our method, we conduct a series of empirical studies on image and molecular datasets, as well as a real-world case study in drug discovery. The results demonstrate that inference set design significantly reduces experimental costs while improving overall system performance. Importantly, this is true even when the generalization benefits of active learning-based acquisition functions are marginal compared to random search. This has important practical implications for active learning: if a problem is a hybrid screen—in the sense that one only needs good performance on a fixed, finite set of experiments—then evaluating generalization error dramatically understates the benefits of active learning. By combining simple active learning acquisition functions with an appropriate stopping criterion, we show that it is possible to make large scale screening far more efficient.
@@ 2 Methods
In this section, we present the problem of efficiently evaluating a large, finite set of experimental designs, such as screening a compound library. We formalize it as a subset selection problem, and show that it can be solved using an active learning strategy. Active learning is effective in this setting because it essentially selects its own inference set; we call this mechanism, inference set design. We propose a stopping criterion to monitor the model’s performance and trigger termination of the acquisition process and show that this algorithm allows to reach provably high levels of performance on a target set of samples. A pseudo-code is available in Appendix B.
@@ 2.1 Reducing Experimental Costs via Hybrid Screens and Inference Set Design
We are motivated by the problem of efficient data acquisition of compound libraries for drug discovery applications. Each “experimental readout” (or label) $y$ requires us to run an experiment, which has some associated cost. We assume that we have a predefined target set $\mathcal{X}_{\text{target}}=\{x_{i}\}^{N_{\text{target}}}_{i=1}$ of experimental designs for which we want to acquire a readout, where $N_{\text{target}}$ represents the number of samples we wish to obtain labels or predictions for. For example, in drug discovery, a common set of experimental designs, $\mathcal{X}_{\text{target}}$, would be a library of thousands or even millions of compounds, each required to be tested on several cell types and at different concentrations. For each $x_{i}\in\mathcal{X}_{\text{target}}$, the associated experimental readout, $y_{i}$, would correspond to the drug’s effect on the corresponding cell type. One way to reduce the acquisition costs of screening the entire target set is to train a predictive model $\hat{p}(x)$ on the observation set – the subset of the target set, $\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}_{\text{target}}$, for which we already have observed the labels $y$. We can then use this model’s predictions in place of the real labels on the remaining samples. Because experiments are typically run in multiple rounds, at each round $t$, the target set $\mathcal{X}_{\text{target}}$ can be partitioned into two mutually exclusive subsets:
@@
  • The observation set $\mathcal{X}_{\text{obs}}^{t}$, which contains the union of all experimental designs that have already been tested and for which the readouts, $\mathcal{Y}_{\text{obs}}^{t}$, have been observed. We denote their pairing the observation dataset $\mathcal{D}_{\text{obs}}^{t}:=(\mathcal{X},\mathcal{Y})_{\text{obs}}^{t}:=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{obs}}}$. At each step $t$, a constant number of $N_{b}$ samples are acquired and added to the observation set, causing this set to grow over time.
@@
  • The inference set $\mathcal{X}_{\text{inf}}^{t}:=\mathcal{X}_{\text{target}}\setminus\mathcal{X}_{\text{obs}}^{t}$, which consists of the remaining experimental designs in the target set that have not yet been tested. This set shrinks over time as samples are selected from it for acquisition and transferred into the observation set.
@@
Once acquisition is stopped, at timestep $t=\tau$, the output of the system is a hybrid dataset, i.e. the combination $(\mathcal{X},\mathcal{Y})_{\text{obs}}^{\tau}\cup(\mathcal{X},\hat{\mathcal{Y}})^{\tau}_{\text{inf}}$ of both the labeled pairs from the observation set and the predicted pairs from the inference set. We call this procedure a hybrid screen, as it provides a hybrid set of labels for the entire target set – some measured, some predicted. The system’s performance at each step $t$ is evaluated based on its accuracy $\mu_{\text{sys}}^{t}$ across the entire target set, which consists of perfectly accurate labels $y$ from the observation set and the inferred labels $\hat{y}$ for the inference set: [^1] Equation 1 assumes a classification problem, but we also explore continuous outcomes $y$ in Appendix E.1.
@@
where $|\mathcal{X}_{\text{obs}}|$ denotes the number of samples in the observation set. A natural approach is to treat this as an active learning problem. Like standard active learning, we can sequentially select subsets of data for labeling and the system accuracy will benefit from any improvements in generalization that arise from better selection of training examples. However, this setting has an important difference: we only ever evaluate $\hat{p}$ on $\mathcal{X}_{\text{inf}}$, and this set is under the control of the learner because any example that is selected for labeling is removed from the inference set. As a result, by choosing acquisition functions that select harder training examples first, we make the test-time task of the learner easier, as the harder examples are effectively pruned out of the inference set. We call this approach inference set design – by selective label acquisition, the agent actively designs the composition of the inference set such that it can excel on it. This mechanism is illustrated in Figure 1.
@@ 2.2 Empirical Stopping Criterion
The key technical challenge of inference set design is deciding when to stop running new experiments. In a deployment environment, the labels for the inference set will remain unavailable, so we need to estimate the system performance from the collected data only. To avoid unnecessary experimentation, inference set design algorithms must employ an efficient stopping criterion. We address this by maintaining a probabilistic lower bound on the system accuracy, and stopping once this bound exceeds the critical threshold, $\gamma$. A simple approach to maintaining such a bound is to leverage the feedback from each round of experimentation.
@@
$x_{\text{selected}}:=\operatorname*{arg\,max}_{x_{i}\in\mathcal{X}_{\text{inf}}\setminus\mathcal{X}_{b}}\Big{(}1-\hat{v}_{i}\Big{)}\,\,,\quad\text{for}\quad i=1,\dots,N_{b}$ (3)
@@ 3 Related Work
In this work, we apply Active Learning (AL) to the problem of hybrid screens for biological data acquisition. Warmuth et al. (2003) were among the first to apply AL to the field of drug discovery, with an acquisition function based on support vector machines for binding affinity predictors. AL has since then been explored as a tool for problems such as virtual screening (Fujiwara et al., 2008), cancer monitoring (Danziger et al., 2009), protein-protein interaction (Mohamed et al., 2010), compound classification (Lang et al., 2016) and chemical space exploration (Smith et al., 2018).
These approaches seek to produce a better model than training on random samples would, and are to be distinguished from bayesian optimization methods (Graff et al., 2021; Gorantla et al., 2024), which also fall under the umbrella of “active learning” but instead use the model to seek specific samples (e.g. with high binding affinity) and often focus on addressing the exploration-exploitation dilemma which has been widely studied by bandit algorithms (Svensson et al., 2022). For hybrid screens, we seek a model that can accurately predict the labels of all samples in the inference set, without attempting to maximize the value of the readouts. Our setting, focusing on a fixed set of data, is sometimes referred to as pool-based active learning (Wu, 2019; Zhan et al., 2021). A key difference is that most active learning work (including pool-based AL) tend to focus evaluations on a held-out test set (Atlas et al., 1989; Cohn et al., 1994; Gal et al., 2017; Margatina et al., 2021; Zhan et al., 2021; Luo et al., 2023; Li et al., 2024) or restrict the algorithm to very limited acquisition budgets (e.g. only $10\%$ of the data pool) (Wu, 2019), which generally does not leave enough margin for mechanisms such as inference set design to dominate.
@@ 4 Experiments
In our experiments [^2] The code is available at https://github.com/ineporozhnii/inference_set_design. All datasets to reproduce our results are publicly available, except one proprietary dataset for the results in Figure 8., we aim at highlighting the key role that inference set design plays in allowing AL agents to reach high levels of performance on the inference set (Section 4.1), validating the robustness of this mechanism on molecular datasets used for drug discovery (Section 4.2), applying this method to a challenging real-world case of compound library screening on a large scale phenotypic assays (Section 4.3) and empirically validating our theoretical assumption and stopping criterion (Section 4.4). Unless specified otherwise, all curves show average metrics across 3 seeds and the shaded areas denote the standard error of the mean. Vertical lines denote the average stopping time, with min and max intervals (left to right). For all modalities, we use vectorized representations as input and our model $\hat{p}$ is parameterized using a feed-forward deep neural network with residual connections. Additional details regarding data pre-processing, network architectures and hyperparameters are presented in Appendices C and D.
Figure 2: Performance of least-confidence (LC) and random acquisition functions on variations of MNIST (columns) across 3 seeds. a) Original: MNIST with its naturally occurring easy and hard examples. b) Partial Observability: MNIST where the bottom two thirds of the images have been cropped out. c) Noisy labeling function: MNIST with shuffled labels for images of 6’s, 8’s, and 9’s. The system seeks a target accuracy $\gamma$ of $98\%$.
Figure 3: Performance of active and random agents on MNIST subject to an increasing number of shuffled labels. As the task becomes harder, the accuracy gap increases, but the stopping time $\tau$ is triggered later.